# Rethinking Multiple Instance Learning for Whole Slide Image Classification: A Good Instance Classifier is All You Need

### v1

@@ Abstract
Weakly supervised whole slide image classification is usually formulated as a multiple instance learning (MIL) problem, where each slide is treated as a bag, and the patches cut out of it are treated as instances. Existing methods either train an instance classifier through pseudo-labeling or aggregate instance features into a bag feature through attention mechanisms and then train a bag classifier, where the attention scores can be used for instance-level classification. However, the pseudo instance labels constructed by the former usually contain a lot of noise, and the attention scores constructed by the latter are not accurate enough, both of which affect their performance. In this paper, we propose an instance-level MIL framework based on contrastive learning and prototype learning to effectively accomplish both instance classification and bag classification tasks. To this end, we propose an instance-level weakly supervised contrastive learning algorithm for the first time under the MIL setting to effectively learn instance feature representation. We also propose an accurate pseudo label generation method through prototype learning. We then develop a joint training strategy for weakly supervised contrastive learning, prototype learning, and instance classifier training. Extensive experiments and visualizations on four datasets demonstrate the powerful performance of our method. Codes will be available.
{IEEEkeywords}
@@
1 Introduction
The deep learning-based whole slide image (WSI) processing technology is expected to greatly promote the automation of pathological image diagnosis and analysis [1, 2, 3, 4, 5]. However, WSIs are quite different from natural images in size, and the size can range from 100 million to 10 billion pixels, which makes it impossible to directly utilize deep learning models developed for natural images to WSIs. It is a common approach to divide WSI into many non-overlapping small patches for processing, but providing fine-grained annotations for these patches is very expensive (a WSI can typically produce tens of thousands of patches), making patch-based supervised methods infeasible [6, 7, 8, 9, 10]. Therefore, weakly supervised learning approaches based on Multiple Instance Learning (MIL) have become the mainstream in this field. In the MIL setting, each WSI is regarded as a bag, and the small patches cut out of it are regarded as instances of the bag. For a positive bag, there is at least one positive instance, while for a negative bag, all instances are negative. In clinical applications, there are two main tasks for WSI classification: bag-level classification, which accurately predicts the class of a whole slide, and instance-level classification, which accurately identifies positive instances [11, 12].
Figure 1: Motivation of our method. A. Existing instance-based MIL methods typically assign a bag’s label to its instances as pseudo labels, resulting in a large number of noises in instance pseudo labels. B. The loss function of bag-based MIL methods is defined at the bag level, which often only finds the most easily identifiable positive instances and ignores other more difficult ones. C. In contrast to these methods, we propose an effective instance-based MIL framework based on contrastive learning and prototype learning and a joint training strategy.
@@ 1 Introduction
Bag-based methods first use an instance-level feature extractor to extract features for each instance in a bag, and then aggregate these features to obtain a bag-level feature, which is used to train a bag classifier. Most recent bag-based methods utilize attention mechanisms to aggregate instance features and they introduce an independent scoring module to generate learnable attention weights for each instance feature, which can be used to realize instance-level classification. Although this type of method overcomes the problem of noisy labels in instance-based methods, it also has the following issues: 1) Low performance in instance-level classification. We find that the difficulty of identifying different positive instances is different in the same positive bag (e.g., instances with larger tumor areas are easier to be identified than those with smaller tumor areas). Attention-based methods define losses at the bag level, which often leads to the result that only the most easily identifiable positive instances are found while other more difficult ones are missed [19, 12]. In other words, the network assigns higher attention weights to the most easily recognizable positive instances to achieve correct bag classification, without the need to find all positive instances, which greatly limits its instance-level classification performance. 2) Bag-level classification performance is not robust. As mentioned earlier, bag-level classification relies heavily on the attention scores assigned by the scoring network to each instance. When these attention scores are inaccurate, the performance of the bag classifier will also be affected. A typical example is the bias that occurs in classifying bags with a large number of difficult positive instances while very few easy positive instances. Figure 1 B shows the typical training process and issues of bag-based methods.
@@
In the history of deep learning-based MIL research, although instance-based methods were first proposed, their reliance on instance pseudo labels, which are difficult to obtain, has led to a bottleneck in their performance. In contrast, an increasing number of researchers [20, 22] have focused on using stronger attention-based methods at the bag level to provide more accurate attention scores. However, intuitively, as long as the loss function is defined at the bag level, attention-based scoring methods will inevitably exhibit ”laziness” in finding more difficult positive instances [19, 12]. Different from the above-mentioned studies, we propose an instance-based MIL framework based on contrastive learning and prototype learning, called INS. Figure 1 C illustrates the main idea and basic components of INS. Our main objective is to directly train an efficient instance classifier at the fine-grained instance level, which needs to fulfill two requirements: first, obtaining a good instance-level feature representation, and second, assigning an accurate pseudo label to each instance. To this end, we propose instance-level weakly supervised contrastive learning (IWSCL) for the first time in the MIL setting to learn good instance representations, better separating negative and positive instances in the feature space. We also propose the Prototype-based Pseudo Label Generation (PPLG) strategy, which generates high-quality pseudo labels for each instance by maintaining two representative feature vectors as prototypes, one for negative instances and the other for positive instances. We further develop a joint training strategy for IWSCL, PPLG and the instance classifier. Overall, IWSCL and PPLG are completed under the guidance of the instance classifier’s predictions. At the same time, the good feature representations from IWSCL and the accurate pseudo labels from PPLG can then further improve the instance classifier. More importantly, we efficiently utilize the true negative instances from negative bags in the training set to guide all the instance classifier, the IWSCL and the PPLG, ensuring that INS iterates towards the right direction. After training the instance classifier, due to its strong instance classification performance, we can complete accurate bag classification using simple mean pooling.
@@
We comprehensively evaluated the performance of INS on six tasks, using a simulated CIFAR10 dataset and three real-world datasets containing breast cancer, lung cancer, and cervical cancer. Extensive experimental results demonstrate that INS achieved better performance in both instance and bag classification than state-of-the-art methods. More importantly, our experiments not only include the tasks that human doctors can directly judge from H&\&&E-stained slides, like tumor diagnosis and tumor subtyping, but also the tasks that human doctors cannot directly make decisions from HE slides, including predicting lymph node metastasis from primary lesion, patient prognosis, and prediction of immunohistochemical markers. Given the strong instance-level classification ability of INS, we can use it for explainable research and new knowledge discovery in these difficult clinical tasks. In the task of predicting lymph node metastasis from primary lesion of cervical cancer, we use INS to classify high- and low-risk instances, thereby identifying the ”Micropapillae” pathological pattern that indicates high risk of lymph node metastasis.
@@
2 Related Work
2.1 Instance-based MIL Methods
@@ 2.2 Bag-based MIL Methods
Bag-based methods first extract instance features and then aggregate these features in a bag to generate bag features for training a bag classifier. Attention-based methods [16, 17, 18, 19, 20, 21, 10] are the mainstream of this paradigm, which typically use an independent scoring network for each instance feature to produce learnable attention weights, which can also be used to generate instance predictions. The main problem of these methods is that they cannot accurately identify difficult positive instances, resulting in limited instance and bag classification performance. Recently, some studies [19, 12, 24] have added instance-level classification loss to bag-level losses, but the pseudo labels assigned to instances are still noisy. Some methods have also been proposed to accomplish WSI classification using reinforcement learning [30], graph learning [31, 32], and bayesian learning [33]. However, none of them can effectively fulfill the instance classification task. In contrast, we directly start from the finest instance-level and use weakly supervised contrastive learning and prototype learning to complete instance feature learning and pseudo-label updating, thereby addressing both instance and bag classification.
@@
Figure 2: Comparison of our IWSCL with existing contrastive learning methods. The red box and blue box represent positive and negative instances, respectively. ”Ins” is the abbreviation for instances, ”pos” for positive, ”neg” for negative, and ”CL” for contrastive learning. A. Self-supervised instance-level contrastive learning separates instance features as much as possible, regardless of whether they are positive or negative. B. Weakly supervised contrastive learning at the bag level [25] does not take instance features into account. C. The proposed instance-level weakly supervised contrastive learning tries to separate positive and negative instances directly in the feature space.
2.3 Contrastive Learning for WSI Classification
@@
Existing methods [20, 21, 22, 12, 34, 35, 36] usually first use WSI patches to pretrain an instance-level feature extractor through self-supervised learning and then perform model training using the extracted features. Most of them [20, 21, 12, 35, 36] use contrastive self-supervised learning methods [37, 38, 39] to extract instance features, but this process is completely unsupervised and it can only attempt to separate all instances as much as possible instead of effectively separating positive and negative instances. In the latest research, Wang et al. [25] proposed a feature-based contrastive learning method at the bag-level, but it still cannot perform effective instance classification. In contrast, we for the first time propose instance-level weakly supervised contrastive learning (IWSCL) under the MIL setting, which effectively separates negative and positive instance features. Figure 2 shows the differences between existing contrastive learning methods and our proposed IWSCL.
2.4 Prototype Learning for WSI Classification
@@
Prototype learning, derived from Nearest Mean Classifiers, aims to provide a concise representation for instances. Recent studies demonstrate the potential of using representations or prototypes for classification, with variations in construction and utilization. PMIL [40] employs unsupervised clustering to construct prototypes, enhancing bag features by assessing instance similarity within bags. TPMIL [41] creates learnable prototype vectors, utilizing attention scores as soft pseudo-labels to assign instances. However, PMIL’s non-learnable prototypes focus on improving bag features, posing challenges for fine-grained instance classification. TPMIL heavily relies on attention scores, which often fail to accurately identify challenging positive instances. Additionally, their prototype learning lacks effective integration of feature-level contrastive learning, resulting in limited performance. In contrast, our proposed PPLG strategy generates high-quality pseudo-labels, utilizing joint training with instance contrastive representation learning, prototype learning, and the instance classifier. Our prototype learning also incorporates guidance from the instance classifier and includes true negative instances from negative bags in the training set.
3 Method
3.1 Problem Formulation
@@ 3.2 Framework Overview
Figure 3: Workflow of the proposed INS framework, where ”pos” is the abbreviation for positive, ”neg” for negative, ”Aug” for augmentation, ”IWSCL” for the proposed instance-level weakly supervised contrastive learning, and ”PPLG” for the proposed prototype-based pseudo label generation.
Specifically, in one iteration, we first randomly select an instance $x_{i,j}$ from all instances in the training set, and generate a query view and a key view using two different augmentations. In the Query View branch, we input the query view into an Encoder and then feed its output to both the instance classifier and the MLP-based projector to obtain the predicted class ${\hat{y}}_{i,j}\in\mathbb{R}^{2}$ (a one-hot vector indicating negative or positive) and the feature embedding $q_{i,j}\in\mathbb{R}^{d}$, respectively. In the Key View branch, we input the key view into an Encoder and then feed its output to a Projector to obtain the feature embedding $k_{i,j}\in\mathbb{R}^{d}$, where both Encoder and Projector are updated through momentum-based methods from the Query View branch. Inspired by MOCO [38] and Pico [42], we maintain a large Embedding Queue to store the feature embeddings of the Key View branch together with the predicted class labels of the corresponding instances. Then, we use the current instance’s ${\hat{y}}_{i,j}$$q_{i,j}$, $k_{i,j}$, and the Embedding Queue from the previous iteration to perform instance-level weakly supervised contrastive learning (IWSCL). For $q_{i,j}$, we pull closer the instance embeddings in the Embedding Queue that have the same predicted class and push away those with different predicted class. In the PPLG module, we maintain two representative feature vectors as prototypes for positive and negative classes during training. We use $q_{i,j}$ and the prototype vectors from the previous iteration to generate pseudo labels for $x_{i,j}$. Finally, we update the Embedding Queue using ${\hat{y}}_{i,j}$ and $k_{i,j}$, update the prototype vectors using ${\hat{y}}_{i,j}$ and $q_{i,j}$, and train the instance classifier with the generated pseudo label, which completes the current iteration. In addition, to prevent bag-level degradation during training, we add a bag-level constraint loss function. The IWSCL module and the PPLG module are presented in Section 3.3 and Section 3.4, respectively. The bag constraint and total loss are given in Section 3.5.
3.3 Instance-level Weakly Supervised Contrastive Learning
@@
3.4 Prototype-based Pseudo Label Generation
@@
3.5 Bag Constraint and Total Loss
@@
4 Experimental Results
4.1 Datasets

### v2

@@ Abstract
Weakly supervised whole slide image classification is usually formulated as a multiple instance learning (MIL) problem, where each slide is treated as a bag, and the patches cut out of it are treated as instances. Existing methods either train an instance classifier through pseudo-labeling or aggregate instance features into a bag feature through attention mechanisms and then train a bag classifier, where the attention scores can be used for instance-level classification. However, the pseudo instance labels constructed by the former usually contain a lot of noise, and the attention scores constructed by the latter are not accurate enough, both of which affect their performance. In this paper, we propose an instance-level MIL framework based on contrastive learning and prototype learning to effectively accomplish both instance classification and bag classification tasks. To this end, we propose an instance-level weakly supervised contrastive learning algorithm for the first time under the MIL setting to effectively learn instance feature representation. We also propose an accurate pseudo label generation method through prototype learning. We then develop a joint training strategy for weakly supervised contrastive learning, prototype learning, and instance classifier training. Extensive experiments and visualizations on four datasets demonstrate the powerful performance of our method. Codes are available at https://github.com/miccaiif/INS.
Index Terms:
@@
I Introduction
Figure 1: Motivation of our method. (A). Existing instance-based MIL methods typically assign a bag’s label to its instances as pseudo labels, resulting in a large number of noises in instance pseudo labels. (B). The loss function of bag-based MIL methods is defined at the bag level, which often only finds the most easily identifiable positive instances and ignores other more difficult ones. (C). In contrast to these methods, we propose an effective instance-based MIL framework based on contrastive learning and prototype learning and a joint training strategy. Our main goal is to develop an effective instance classifier using instance-level weakly supervised contrastive learning and pseudo-label generation. Contrastive learning aims to improve instance representations to better distinguish between negative and positive instances in the feature space. High-quality pseudo-labels are then generated for each instance based on representative feature vectors. Both feature learning and pseudo-label generation are guided by the instance classifier, enhancing its capabilities.
The deep learning-based whole slide image (WSI) processing technology is expected to greatly promote the automation of pathological image diagnosis and analysis [1, 2, 3, 4, 5, 6, 7]. However, WSIs are quite different from natural images in size, and the size can range from 100 million to 10 billion pixels, which makes it impossible to directly utilize deep learning models developed for natural images to WSIs. It is a common approach to divide WSI into many non-overlapping small patches for processing, but providing fine-grained annotations for these patches is very expensive (a WSI can typically produce tens of thousands of patches), making patch-based supervised methods infeasible [8, 9, 10, 11, 12]. Therefore, weakly supervised learning approaches based on Multiple Instance Learning (MIL) have become the mainstream in this field. In the MIL setting, each WSI is regarded as a bag, and the small patches cut out of it are regarded as instances of the bag. For a positive bag, there is at least one positive instance, while for a negative bag, all instances are negative. In clinical applications, there are two main tasks for WSI classification: bag-level classification, which accurately predicts the class of a whole slide, and instance-level classification, which accurately identifies positive instances [13, 14].
@@ I Introduction
Bag-based methods first use an instance-level feature extractor to extract features for each instance in a bag, and then aggregate these features to obtain a bag-level feature, which is used to train a bag classifier. Most recent bag-based methods utilize attention mechanisms to aggregate instance features and they introduce an independent scoring module to generate learnable attention weights for each instance feature, which can be used to realize instance-level classification. Although this type of method overcomes the problem of noisy labels in instance-based methods, it also has the following issues: 1) Low performance in instance-level classification. We find that the difficulty of identifying different positive instances is different in the same positive bag (e.g., instances with larger tumor areas are easier to be identified than those with smaller tumor areas). Attention-based methods define losses at the bag level, which often leads to the result that only the most easily identifiable positive instances are found through the high attention scores while other more difficult ones are missed [21, 14]. In other words, the network only assigns higher attention weights to the most easily recognizable positive instances to achieve correct bag classification, without the need to find all positive instances, which greatly limits its instance-level classification performance. 2) Bag-level classification performance is not robust. As mentioned earlier, bag-level classification relies heavily on the attention scores assigned by the scoring network to each instance. When these attention scores are inaccurate, the performance of the bag classifier will also be affected. A typical example is the bias that occurs in classifying bags with a large number of difficult positive instances while very few easy positive instances. Figure 1 B shows the typical training process and issues of bag-based methods.
@@
In the history of deep learning-based MIL research, although instance-based methods were first proposed, their reliance on instance pseudo labels, which are difficult to obtain, has led to a bottleneck in their performance. In contrast, an increasing number of researchers [22, 24] have focused on using stronger attention-based methods at the bag level to provide more accurate attention scores. However, intuitively, as long as the loss function is defined at the bag level, attention-based scoring methods will inevitably exhibit "laziness" in finding more difficult positive instances [21, 14]. Different from the above-mentioned studies, we propose an instance-based MIL framework based on contrastive learning and prototype learning, called INS. Figure 1 C illustrates the main idea and basic components of INS. Our main objective is to directly train an efficient instance classifier at the fine-grained instance level, which needs to fulfill two requirements: first, obtaining a good instance-level feature representation, and second, assigning an accurate pseudo label to each instance. To this end, we propose instance-level weakly supervised contrastive learning (IWSCL) for the first time in the MIL setting to learn good instance representations, better separating negative and positive instances in the feature space. We also propose the Prototype-based Pseudo Label Generation (PPLG) strategy, which generates high-quality pseudo labels for each instance by maintaining two representative feature vectors as prototypes, one for negative instances and the other for positive instances. We further develop a joint training strategy for IWSCL, PPLG and the instance classifier. Overall, IWSCL and PPLG are completed under the guidance of the instance classifier’s predictions. At the same time, the good feature representations from IWSCL and the accurate pseudo labels from PPLG can then further improve the instance classifier. More importantly, we efficiently utilize the true negative instances from negative bags in the training set to guide all the instance classifier, the IWSCL and the PPLG, ensuring that INS iterates towards the right direction. After training the instance classifier, due to its strong instance classification performance, we can complete accurate bag classification using simple mean pooling.
@@
We comprehensively evaluated the performance of INS on six tasks, using a simulated CIFAR10 dataset and three real-world datasets containing breast cancer, lung cancer, and cervical cancer. Extensive experimental results demonstrate that INS achieved better performance in both instance and bag classification than state-of-the-art methods. More importantly, our experiments not only include the tasks that human doctors can directly judge from H&\&&E-stained slides, like tumor diagnosis and tumor subtyping, but also the tasks that human doctors cannot directly make decisions from HE slides, including predicting lymph node metastasis from primary lesion, patient prognosis, and prediction of immunohistochemical markers. Given the strong instance-level classification ability of INS, we can use it for explainable research and new knowledge discovery in these difficult clinical tasks. In the task of predicting lymph node metastasis from primary lesion of cervical cancer, we use INS to classify high- and low-risk instances, thereby identifying the "Micropapillae" pathological pattern that indicates high risk of lymph node metastasis.
@@
II Related Work
II-A Instance-based MIL Methods
@@ II-B Bag-based MIL Methods
Bag-based methods first extract instance features and then aggregate these features in a bag to generate bag features for training a bag classifier. Attention-based methods [18, 19, 20, 21, 22, 23, 12, 32] are the mainstream of this paradigm, which typically use an independent scoring network for each instance feature to produce learnable attention weights, which can also be used to generate instance predictions. The main problem of these methods is that they cannot accurately identify difficult positive instances, resulting in limited instance and bag classification performance. Recently, some studies [21, 14, 26] have added instance-level classification loss to bag-level losses, but the pseudo labels assigned to instances are still noisy. Some methods have also been proposed to accomplish WSI classification using reinforcement learning [33], graph learning [34, 35], and bayesian learning [36]. However, none of them can effectively fulfill the instance classification task. In contrast, we directly start from the finest instance-level and use weakly supervised contrastive learning and prototype learning to complete instance feature learning and pseudo-label updating, thereby addressing both instance and bag classification.
Figure 2: Comparison of our instance-level weakly supervised contrastive learning (IWSCL)
@@
with existing contrastive learning methods. The red box and blue box represent positive and negative instances, respectively. "Ins" is the abbreviation for instances, "pos" for positive, "neg" for negative, and "CL" for contrastive learning. A. Self-supervised instance-level contrastive learning separates instance features as much as possible, regardless of whether they are positive or negative. B. Weakly supervised contrastive learning at the bag level [27] does not take instance features into account. C. The proposed IWSCL tries to separate positive and negative instances directly in the feature space.
II-C Contrastive Learning for WSI Classification
@@
Existing methods [22, 23, 24, 14, 37, 38, 39] usually first use WSI patches to pretrain an instance-level feature extractor through self-supervised learning and then perform model training using the extracted features. Most of them [22, 23, 14, 38, 39] use contrastive self-supervised learning methods [40, 41, 42, 43, 44, 45, 46] to extract instance features, but this process is completely unsupervised and it can only attempt to separate all instances as much as possible instead of effectively separating positive and negative instances. In the latest research, Wang et al. [27] proposed a feature-based contrastive learning method at the bag-level, but it still cannot perform effective instance classification. In contrast, we for the first time propose instance-level weakly supervised contrastive learning (IWSCL) under the MIL setting, which effectively separates negative and positive instance features. Figure 2 shows the differences between existing contrastive learning methods and our proposed IWSCL.
II-D Prototype Learning for WSI Classification
@@
Prototype learning, derived from Nearest Mean Classifiers, aims to provide a concise representation for instances. Recent studies demonstrate the potential of using representations or prototypes for classification, with variations in construction and utilization [47, 48, 49, 50]. ProtoMIL [51] initially extracts features from each patch using a convolutional layer, then builds negative and positive prototypes by comparing them through similarity calculations, and finally aggregates them using attention scores. PMIL [52] employs unsupervised clustering to construct prototypes, enhancing bag features by assessing instance similarity within bags. TPMIL [53] creates learnable prototype vectors, utilizing attention scores as soft pseudo-labels to assign instances. However, PMIL’s non-learnable prototypes focus on improving bag features, posing challenges for fine-grained instance classification. ProtoMIL and TPMIL heavily relies on attention scores, which often fail to accurately identify challenging positive instances. Additionally, their prototype learning lacks effective integration of feature-level contrastive learning, resulting in limited performance. In contrast, our proposed PPLG strategy generates high-quality pseudo-labels, utilizing joint training with instance contrastive representation learning, prototype learning, and the instance classifier. Our prototype learning also incorporates guidance from the instance classifier and includes true negative instances from negative bags in the training set.
III Method
III-A Problem Formulation
@@ III-B Framework Overview
Figure 3: Workflow of the proposed INS framework, where "pos" is the abbreviation for positive, "neg" for negative, "Aug" for augmentation, "IWSCL" for the proposed instance-level weakly supervised contrastive learning, and "PPLG" for the proposed prototype-based pseudo label generation.
Specifically, in one iteration, we first randomly select an instance $x_{i,j}$ from all instances in the training set, and generate a query view and a key view using two different augmentations. In the Query View branch, we input the query view into an Encoder and then feed its output to both the instance classifier and the MLP-based projector to obtain the predicted class ${\hat{y}}_{i,j}\in\mathbb{R}^{2}$ (a one-hot vector indicating negative or positive) and the feature embedding $q_{i,j}\in\mathbb{R}^{d}$, respectively. In the Key View branch, we input the key view into an Encoder and then feed its output to a Projector to obtain the feature embedding $k_{i,j}\in\mathbb{R}^{d}$, where both Encoder and Projector are updated through momentum-based methods from the Query View branch. Inspired by MOCO [41] and Pico [54], we maintain a large Embedding Queue to store the feature embeddings of the Key View branch together with the predicted class labels of the corresponding instances. Then, we use the current instance’s ${\hat{y}}_{i,j}$$q_{i,j}$, $k_{i,j}$, and the Embedding Queue from the previous iteration to perform instance-level weakly supervised contrastive learning (IWSCL). For $q_{i,j}$, we pull closer the instance embeddings in the Embedding Queue that have the same predicted class and push away those with different predicted class. In the PPLG module, we maintain two representative feature vectors as prototypes for positive and negative classes during training. We use $q_{i,j}$ and the prototype vectors from the previous iteration to generate pseudo labels for $x_{i,j}$. Finally, we update the Embedding Queue using ${\hat{y}}_{i,j}$ and $k_{i,j}$, update the prototype vectors using ${\hat{y}}_{i,j}$ and $q_{i,j}$, and train the instance classifier with the generated pseudo label, which completes the current iteration. In addition, to prevent bag-level degradation during training, we add a bag-level constraint loss function. The IWSCL module and the PPLG module are presented in Section III-C and Section III-D, respectively. The bag constraint and total loss are given in Section III-E.
III-C Instance-level Weakly Supervised Contrastive Learning
@@
III-D Prototype-based Pseudo Label Generation
@@
III-E Bag Constraint and Total Loss
@@
IV Experimental Results
IV-A Datasets